OpenStudy (fibonaccichick666):

Anyone good at Differential equations? That would not mind me monopolizing their time?

OpenStudy (fibonaccichick666):

I need to find the First integral?<---explain that please. For the exact equation(maybe?) \[\frac{dx}{z+y}=\frac{dy}{x+z}=\frac{dz}{x+y}\]

OpenStudy (fibonaccichick666):

@eliassaab Any thoughts?

OpenStudy (fibonaccichick666):

or @hero ?

OpenStudy (fibonaccichick666):

Any ideas, even if you don't know are more than welcome

OpenStudy (anonymous):

One obvious solution is x=y=z

OpenStudy (anonymous):

Have you considered integrating everything? Assuming \(x,y,z\) are independent of each other.

OpenStudy (fibonaccichick666):

right, but how can we arrive there. And the issue is that would be partials no?

OpenStudy (anonymous):

Another one x=C1; y=C2 ;z = C2

OpenStudy (fibonaccichick666):

ok what if we expressed it as a vector, could we solve it then?

OpenStudy (anonymous):

Consider using the chain rule like \[ \frac{\text{dy}}{\text{dx}}=\frac{\text {dy} \text{dz}}{\text{dx} \text{dz}} \] and try to find a relation between the varialbles

OpenStudy (anonymous):

Sorry, swith the the dx and dz in the denominator

OpenStudy (fibonaccichick666):

it's ok

OpenStudy (anonymous):

\[ \frac{\text{dy}}{\text{dx}}=\frac{\text {dy} \,\text{dz}}{\text{dz}\, \text{dx}} \]

OpenStudy (fibonaccichick666):

hmmm.... do you think I could set it up as \[(x+z)dz=(x+y)dy=(z+y)dz\]

OpenStudy (anonymous):

Yes, something like that

OpenStudy (fibonaccichick666):

oops \[(x+z)dx=(x+y)dy=(z+y)dz\]

OpenStudy (fibonaccichick666):

That legal? I think it should be

OpenStudy (anonymous):

I think so.

OpenStudy (anonymous):

Sorry, I have to go now.

OpenStudy (fibonaccichick666):

it's ok, thank you :)

OpenStudy (fibonaccichick666):

no one at my university can help with this review question...

OpenStudy (fibonaccichick666):

@hartnn can you help?

OpenStudy (fibonaccichick666):

crazy question, eh? @freckles

OpenStudy (freckles):

lol puzzling so far but I will give it some more attention tomorrow I know you are studying for exam so that sucks but its also 12:33 am here :(

OpenStudy (fibonaccichick666):

1:34 here :( But it's alright. I'll just fail haha. The whole class will....

OpenStudy (fibonaccichick666):

yay for curves!

OpenStudy (freckles):

It is always good to never be alone in failing

OpenStudy (fibonaccichick666):

true

OpenStudy (freckles):

or at least it makes you feel a little better about failing

OpenStudy (fibonaccichick666):

also true

OpenStudy (freckles):

http://www.math.utk.edu/~ccollins/M512/HW/soln5.pdf this first question looks like a similar type of equation

OpenStudy (freckles):

Though our equations look like we get know cancellations

OpenStudy (freckles):

like the one there

OpenStudy (freckles):

no*

OpenStudy (freckles):

\[dx(x+z)=dy(z+y) \\ dy(z+y)=dz(x+z) \\ x+z=\frac{dy}{dx}(z+y) \\ dy(z+y)=dz \cdot \frac{dy}{dx}(z+y) \\ (z+y)=dz \frac{z+y}{dx} \\ 1=\frac{dz}{dx}\] make sure I didn't make a mistake

OpenStudy (freckles):

1 dx=1 dz x=z+c

OpenStudy (freckles):

Like so... we can find y now... I suppose

OpenStudy (freckles):

we can replace our x's with (z+c)'s

OpenStudy (fibonaccichick666):

I love you

OpenStudy (fibonaccichick666):

you have two x+z and no y+z

OpenStudy (fibonaccichick666):

\[dx(x+z)=dy(z+y) \\ dy(z+y)=dz(\color\red{x+y}) \\ x+z=\frac{dy}{dx}(z+y) \\ dy(z+y)=dz \cdot \frac{dy}{dx}(z+y) \\ (z+y)=dz \frac{z+y}{dx} \\ 1=\frac{dz}{dx}\]

OpenStudy (freckles):

omg the second line should by (x+y)dy=(z+x)dz right?

OpenStudy (fibonaccichick666):

(x+z)dx=(x+y)dy=(z+y)dz

OpenStudy (fibonaccichick666):

I think it should be like that

OpenStudy (fibonaccichick666):

\[(x+z)dx=(x+y)dy\\(x+y)dy=(z+y)dz\]

OpenStudy (freckles):

goes I keep changing my x's to z' or the other way around

OpenStudy (fibonaccichick666):

I like that better

OpenStudy (freckles):

gosh*

OpenStudy (fibonaccichick666):

haha

OpenStudy (freckles):

let me try to fix this

OpenStudy (freckles):

\[\frac{dx}{z+y}=\frac{dy}{x+z}=\frac{dz}{x+y}\] first to second \[\frac{dx}{z+y}=\frac{dy}{x+z}\] \[(x+z)dx=(z+y)dy\] second to third \[\frac{dy}{z+x}=\frac{dz}{x+y}\] \[(x+y) dy=(z+x)dz\] first to third \[\frac{dx}{z+y}=\frac{dz}{x+y} \\ (x+y) dx=(z+y) dz\]

OpenStudy (fibonaccichick666):

i agree

OpenStudy (freckles):

\[(x+y)dy=(z+x) dz \\ (x+y)dx=(z+y) dz \\ (x+y)=(z+x) \frac{dz}{dy} \text{ first equation \in this mini-post } \\ \text{ plug into second equation \in this mini-post } \\ (z+x) \frac{dz}{dy} dx=(z+y) dz \\ (z+x) \frac{dx}{dy}=z+y\\ \text{ but \in the first equation \in my mini-post \above } \\ \text{ the } (z+x) dx=(z+y) dy \\ \ \\\] this didn't help any because we already had that equation err maybe we can solve those equations for dz above \[\frac{x+y}{z+x} dy=dz \\ \frac{x+y}{z+y} dx=dz \\ \frac{x+y}{z+x} dy=\frac{x+y}{z+y} dx \\ \frac{1}{z+x} dy=\frac{1}{z+y} dx \] still don't know what to do with the z

OpenStudy (fibonaccichick666):

(x+y)dy=(z+x)dz(x+y)dx=(z+y)d< not accurate you mixed up your zs and xs again I think or did something I missed

OpenStudy (freckles):

Yeah I will come back tomorrow sorry

OpenStudy (fibonaccichick666):

it's ok, thanks for trying

OpenStudy (fibonaccichick666):

stay warm and sleep well

OpenStudy (anonymous):

One can show z=x + Constant. let us suppose the z=x, then one get that \[ \frac{\sqrt[3]{\sqrt{e^{6 c_1}-4 e^{3 c_1} x^3}+e^{3 c_1}-2 x^3}}{\sqrt[3]{2}}+\frac{\sqrt[3]{2} x^2}{\sqrt[3]{\sqrt{e^{6 c_1}-4 e^{3 c_1} x^3}+e^{3 c_1}-2 x^3}} \] I got this with Mathematica

OpenStudy (anonymous):

\[ y=\frac{\sqrt[3]{\sqrt{e^{6 c_1}-4 e^{3 c_1} x^3}+e^{3 c_1}-2 x^3}}{\sqrt[3]{2}}+\frac{\sqrt[3]{2} x^2}{\sqrt[3]{\sqrt{e^{6 c_1}-4 e^{3 c_1} x^3}+e^{3 c_1}-2 x^3}} \]